OpenStudy (anonymous):

A tower has a triangular base with side lengths 6y ft, 7y + 2 ft, and 5x – 7 ft. Write a simplified expression for its perimeter.

OpenStudy (anonymous):

6y+7y+2+5x-7 11y+5x-5

OpenStudy (katrinakaif):

6y + 7y + 2 + 5x - 7 Collect like terms such as the "y" terms and constant terms like 2 and - 7 6y + 7y = 13y 2 - 7 = -5 5x cannot add with another term so final answer is 13y + 5x -5

OpenStudy (katrinakaif):

Wait...Eliza how did you get 11y? 6y + 7y = 13 y..?

OpenStudy (anonymous):

oh woops... :) i can't add....

OpenStudy (katrinakaif):

Lol its fine =D

OpenStudy (anonymous):

Medal for everyone.

OpenStudy (anonymous):

:D